# AI-coupled HPC Workflow Applications, Middleware and Performance

### v1

@@ Abstract
AI integration is revolutionizing the landscape of HPC simulations, enhancing the importance, use, and performance of AI-driven HPC workflows. This paper surveys the diverse and rapidly evolving field of AI-driven HPC and provides a common conceptual basis for understanding AI-driven HPC workflows. Specifically, we use insights from different modes of coupling AI into HPC workflows to propose six execution motifs most commonly found in scientific applications. The proposed set of execution motifs is by definition incomplete and evolving. However, they allow us to analyze the primary performance challenges underpinning AI-driven HPC workflows. We close with a listing of open challenges, research issues, and suggested areas of investigation including the the need for specific benchmarks that will help evaluate and improve the execution of AI-driven HPC workflows.
@@ 1 Introduction
Multiple recent publications have demonstrated how AI coupled with traditional HPC simulations can provide practical performance enhancements of 103 or more [1, 2, 3]. AI-coupled-HPC workflows – as opposed to decoupled AI and HPC – involve the concurrent, real-time coupled execution of AI and HPC tasks in ways that allow the AI systems to steer or inform the HPC tasks and vice versa. The real-time coupling and concurrency of HPC and AI tasks are fundamental, as it allows allows bidirectional influence. The online coupling of an AI system to HPC workflows can be used for a variety of scenarios, e.g., an AI-surrogate model to substitute part of expensive simulation, an AI system to learn the function or determine the parameters or the practical fields, and an AI system to guide a campaign using an objective function or performance Pareto-optimal or resource-optimal experiments (using, for example, Active/Reinforcement Learning).
@@
Driving the need to couple AI systems to traditional HPC simulations is the promise of enhancing the “effective performance” of HPC workflows, where enhancement is measured by “science for a given amount of computing”. AI-coupled HPC workflows continue to overcome the limitations of traditional forward simulations in increasingly sophisticated and pervasive ways. Further, AI-coupled HPC workflows will overcome traditional bottlenecks that prevent greater scale – physical and temporal or higher resolutions. The integration of AI into a computational workflow is also a sustainable and scalable way to obtain significant performance gains and present an opportunity to avoid simulation enhancements that are overly sensitive to processor architecture [4, 5]. Put together, AI-coupled HPC workflows present a promising paradigm [6, 7, 8] which leverages the ubiquitous interest and wide capabilities being developed for AI but employs them to overcome performance bottlenecks due to unsustainable approaches.
Integrating AI systems into traditional high-performance computing workflows has demonstrably enabled highly accurate modeling and holds significant promise for accelerating scientific discovery. There exist multiple promising examples of workflows that couple AI methods with traditional HPC workflows, typically by extending HPC workflows with additional capabilities to support the concurrent execution of AI modules  [9, 10, 11, 12]. Multiple solutions have successfully coupled AI systems to be either “about” the HPC workflows or “outside” (typically in the outer loop or even possibly remote from) the main HPC workflow. However, the ability to substitute an AI system “inside” of an HPC simulation remains challenging, with most solutions being customized to specific simulation software or platforms. Figure 1 shows a typical science application where AI has advanced simulations and experiments by coupling AI to HPC workflows via different mechanisms. Furthermore, in molecular science and engineering – which covers a diverse range of domains from biophysics to materials to chemical engineering – AI has been used to make simulation workflows more efficient and effective [13]. AI has been used to predict ever more complex properties, such as energetic materials, solid-state materials properties, and even the structure of complex materials. In addition, several machine-learned potentials that enable access to quantum-chemical-like accuracies at a fraction of the cost have emerged.
@@
The focus in this paper is on the scientific workflows that involve the online (real-time) coupling and concurrent execution of AI and HPC, not offline training and inference. In addition to providing an overview of existing AI-coupled HPC workflows, the primary objectives of this paper are to identify classes of AI-coupled-HPC workflows and to provide an overview of the software systems used to support them. The paper also discusses performance and systems challenges, identifying both robust solutions and the lack thereof. Finally, the paper highlights some open research issues to advance AI-HPC workflows.
The paper is organized as follows: Section 2 reviews coupling modes between AI and HPC tasks and introduces six execution motifs that describe the interaction and coordination patterns of AI and HPC components. These motifs are used to understand a range of AI-driven HPC workflow applications and primary performance challenges underpinning AI-driven HPC workflows. Execution motifs are important at multiple levels, not the least of which is their influence on the design of application-level software systems and task runtime systems. For instance, the frequency of interaction in a given motif informs the performance requirements and the optimizations available to lower-level libraries. Further, the diversity of AI approaches and model types means intrinsic variation in training and execution times. Section 3 presents several frameworks and libraries developed to ease the coupling of AI and HPC components in a single workflow and illustrates how each of these frameworks relates to the proposed high-level modes of execution motifs. Section 4 discusses various performance issues related to AI-HPC workflows – organized around the motifs – such as load balancing, workload management, and dataflow performance bottlenecks. We conclude in Section 5 with a discussion of several open issues and challenges that the field of AI-coupled HPC workflows presents.
@@ 2 Execution Motifs
AI systems can be coupled to HPC workflows using different methods at multiple levels. Such AI-coupled HPC workflows, integrate AI systems with HPC simulations are AI-in-HPC, AI-out-HPC, and AI-about-HPC simulations. These three modes of coupling [6, 5] correspond to the placement of AI methods relative to HPC simulations. AI-in-HPC represents the scenario in which an AI system is introduced instead of a component of the HPC simulation, or possibly, instead of the whole simulation itself, i.e., the AI model serves as a “total surrogate” [14]. AI-out-HPC captures situations where an AI system resides “outside” of the traditional HPC simulation loop but dynamically controls the progression of the HPC workflow [15, 16] (e.g., control of computational campaigns via reinforcement learning). Finally, AI-about-HPC represents the situation where AI systems are concurrent and coupled to the main HPC tasks: analysis and training codes use the output of the HPC simulation to provide further insights. These three modes are not mutually exclusive; in fact, we contend [4, 17, 5] that increasingly, scientific applications will achieve performance gains and methodological advances by combining all three modes. Thus, if not unified, software solutions for managing the execution of each component and their data transfers for all three modes must be compatible.
Building on the concept of coupling modes (AI-x-HPC, where x could be in, out, or about), we introduce execution motifs. Whereas modes provide insight into the static coupling of AI systems with HPC simulations, execution motifs capture and provide information on recurring dynamic interaction patterns between the AI systems and the HPC simulations, driven by a high-level functional goal.
@@
These execution motifs serve as identifiers for (groups of) AI-coupled HPC workflows that exhibit similar interaction and coupling patterns between AI and HPC components, and thus encounter similar performance bottlenecks and require similar optimizations when deployed on HPC systems.
@@
Figure 2: Defining characteristics of AI-coupled HPC workflows: (left) Breakdown of the main characteristics used to define the different behavioral motifs; (right) Main components (boxes) participating in the workflow execution with two or more components interacting given different control and data flow patterns (arrows).
Figure 2 illustrates (left hand-side) one column for each class of characteristics. Figure 2 (right hand-side) shows the main components participating in the workflow execution (namely: traditional HPC simulations and codes; AI analysis and training; instruments; and humans in the loop) together with the interactions between them (dashed lines show the control flow and uninterrupted lines, the data flow). For each motif, different components will be present, and the control and data flow will connect them in different ways. In this section, we will use this diagram to further illustrate the specific interaction and coupling patterns exhibited by each motif.
Based on a literature survey of large-scale workflows, we identify and characterize six primary execution motifs that couple AI and HPC. Specifically, we identify five common execution motifs wherein AI enhances HPC workflows (#1-5), two of them improving both HPC and AI (#3-4) and one where HPC workflows are used to develop or improve AI systems or models (#6). We briefly describe these execution motifs below and summarize their characteristics in Table 1:
1. AI-based steering ensembles of simulations:In this motif, a HPC workflow, such as an ensemble of simulations are “steered” by an AI system. Sometimes, the AI can terminate or spawn simulations, making the coupling between components dynamic. In subsequent tables and figures, we refer to this motif as Steering. 2. Multistage pipeline:This motif represents a campaign composed of multiple stages with a typical AI-based function that reduces/filters simulations to spawn in the next stage based on the output of the current stage. 3. Inverse design:This motif captures the process of calculating, from a set of observations, the causal factors that produced them (e.g., calculating an image in X-ray computed tomography). AI-driven optimizations are used to iteratively identify causal factors from observational data that can be fed back into the experiment or simulation. The AI system can also be improved in the process thanks to the same feedback loop mechanism. 4. Digital replica:This motif includes the concurrent use of simulations and AI digital replicas (e.g., digital twins or multi-fidelity models) for scientific predictions and health monitoring of experiments or high-fidelity simulations. This motif can be executed in real-time, completely autonomously, or involve a scientist in deciding to guide future experiments. 5. Distributed models and dynamic data:This motif corresponds to workflows deployed over wide area networks. These workflows couple HPC simulations and experiments distributed across multiple remote sites with AI-based reduction/analysis and visualization. The AI can be used simply for compression, refactoring, or remote visualization of dynamic data or for steering the simulation or experiments. Workflows in this motif use highly heterogeneous resources, including edge devices, large leadership facilities, local clusters, or individual laptops for running and connecting different components. We refer to this motif as Distributed Models in subsequent tables and figures. 6. Adaptive execution for training large AI models:This motif couples the distributed training of large AI systems with HPC simulations and experiments. Such workflows can include training components using datasets previously generated by HPC simulations or real-time model optimization by running the training sequence concurrent with HPC simulations that are steered to generate the data required to cover the search space more efficiently. In subsequent tables and figures, we refer to this motif as Adaptive Training.
Table 1 summarizes the distinct patterns of each motif according to the characteristics introduced in Figure 2.
Motif / Scope Interaction Patterns Coupling Patterns Example Use Case
Steering
AI improving HPC
- Control and data flow in one direction: data from HPC to AI, control from AI to HPC
- One AI to one or many HPC
- Optionally human in the loop
- Real-time requirements
- Dynamic composition with HPC simulations spawned or terminated on the fly
- Usually running in one facility
AI-out-HPC
@@
- Command-and-control of physical experiments and simulations (e.g. between shots feedback for plasma physics)
Multistage Pipeline
AI improving HPC
- Data flows in one direction from HPC to one or many AI or HPC components
- AI filters control many HPC simulations
- Typically interaction done without human in the loop
- Real-time requirements
- Dynamic composition with branching in the workflow based on filters
- Running in one facility
AI-in-HPC and AI-out-HPC
@@
- Large-scale MD simulations using AI sampling of a system with many degrees of freedom
InverseDesign
AI improving HPC
HPC improving AI
- Control flow from AI to HPC
- Multiple HPC simulations and/or instruments sending data to AI (one or many)
- Typically interaction done without human in the loop
- Real-time is optional (AI can use existing datasets)
- Execution can be concurrent or asynchronous
- Running in one facility
AI-in-HPC
@@
- Materials discovery to address the problem of data sparsity and reduce the need for domain-specific knowledge
Digital Replica
AI improving HPC
HPC improving AI
- Data/control flow in both directions combining experiments and/or HPC with AI
- Multiple HPC to one AI
- Potentially human-in-the-loop to adjust future runs
- Real-time requirements with monitoring and visualization
- Static composition with only visualization started on the fly
- Components can potentially be executed distributed
AI-about-HPC
@@
- Digital twin of a fusion reactor running concurrent with ITER [18] and monitoring
@@
Distributed Models
AI improving HPC
- Data and control flow in both direction (AI to HPC and other AI and HPC to AI)
- One or many HPC and many AI (analysis, replicas)
- Optional human in the loop
- Near real-time requirements
- Typically static composition, dynamic for adapting to the WAN performance (e.g. compression accuracy)
- Geographically distributed
AI-out-HPC and AI-about-HPC
- Edge-to-HPC execution of
large-scale simulations
- Interactive notebook-
based analyses of instrument data and AI output
AdaptiveTraining
HPC improving AI
- Data flows from one HPC simulation to multiple AI
- Control flows from analysis to HPC and AI training
- No human-in-the-loop
- Near real-time requirements for AI, optional with HPC
- Dynamic composition
- Running in one facility, possible federated in the future
AI-out-HPC
- Hyperparameter optimization workflows
@@
AI-based steering ensembles of simulations
@@
(a) AI system steering (ensembles of) high-fidelity HPC simulations or experiments to avoid undesired results or guide the workflow towards better solutions.
(b) Sequential coupling of data-dependent multistage AI-coupled HPC workflow as a pipeline with filtering/analysis codes between stages.
Figure 3: Workflow computational patterns for the Steering (a) and Multistage Pipeline (b) motifs
The following literature survey focuses on workflows for the Steering execution motif. These workflows use AI systems to analyze how “well” (ensembles of) simulations are progressing toward their initial scientific objective. Based on that analysis, they can change the “trajectory” when exploring the parameter space. This allows them to avoid undesired results either or increase performance or accuracy. The training of the AI system is done offline before the execution of the workflow. Workflows in which the steering process is also used to guide the online training of the AI system will be investigated as part of the Adaptive Training motif.
@@
DeepDriveMD [2] is a general-purpose framework to run an analysis code (as a separate application) that pre-processes the simulation data (e.g. selecting only a subset of atoms of interest or calculating physical parameters for a protein’s native state), and then uses machine learning (ML) models on the pre-processed data to decide which simulations to run next and terminates the less productive MD simulations. By focusing on promising regions of the parameter space, AI-based steering can efficiently navigate complex, high-dimensional systems that were previously intractable. AI systems can also significantly reduce the computational resources required to obtain accurate results by steering simulations toward more relevant configurations.
The aforementioned example is a typical implementation of the Steering motif, but other examples using different coupling solutions while sharing the characteristics presented in Table 1 exist. The AI system analyzes the output of the HPC simulation and either steers the existing simulation or triggers new simulations based on the outcome.
@@
High-Throughput Virtual Screening Pipelines are emblematic of this motif. For instance, a virtual drug discovery pipeline can use AI to improve the effective sampling of individual stages [25]. Integrating multiple workflows, which can each be an AI-coupled HPC workflow, into a single multistage pipeline is another example of AI-driven discovery [26]. AI can also be used in workflows composing multiple multi-fidelity simulations to dynamically adjust how many and what fidelity levels to run based on the desired outcome [27].
Less complex multistage pipelines can be found in neuroscience [28, 29] and cancer research [30]. Both fields employ workflows composed of a sequence of medical image processing applications that process the same Magnetic Resolution Image (MRI) or Whole Slide Image (WSI) with the output of one application informing the following decision. Each image-processing application is typically a mix of HPC and AI codes composed of multiple stages, including cleaning, standardizing, analysis, and post-processing within one large code or as separate applications. One such workflow is, for instance, used to identify lesions in the brain [29]. It consists of a 3-stage pipeline, starting with a set of MRI images that go through a whole segmentation code, followed by a surface reconstruction application, and finally by an application to create a structure of the brain in MRI space. Based on the output of these three applications, the workflow can further trigger additional codes depending on the research objectives.
@@
Deep learning has been used to solve inverse problems, i.e., determining causal factors from a set of observations, in many scientific fields, including material science [33, 34], nanophotonics [35, 36, 37], or heat transfer [38]. However, the lack of sufficiently large training data sets is often a crucial issue for these learning algorithms since their accuracy heavily relies on the amount and the quality of the available data. This is, for instance, the case with material characteristics identification, which requires a huge amount of user time to acquire observational data. HPC simulations have recently started generating sufficient noise-free data to train AI models. The new challenge is getting accurate simulation results, often requiring extensive testing and expertise to fine-tune input parameters. This can also be time-consuming and resource-intensive, especially for complex simulations since it requires validation over many tests. AI-coupled HPC workflows following the Inverse Design motif address this data sparsity problem and reduce the demand for domain-specific knowledge in determining the input parameters of a simulation.
@@
(a) Iterative refinement of design parameters combining experiments with HPC simulations and AI surrogates
(b) Physics-based models combining HPC high-fidelity simulations and AI reduced order models and/or digital twins
@@
Fusion energy researchers are turning to digital twins as a promising avenue for efficient whole-device optimization and virtual component qualification, crucial for successfully designing fusion reactor components under conditions that defy real-world testing. Digital twin workflows are primed to play a crucial role in ensuring the success of ITER [18], and pave the way for the future of commercial fusion energy. Several studies [49, 50, 51] presents the vision of the fusion science community to develop future workflows. They all discuss the design of a workflow combining digital twins of different parts of the fusion reactor (e.g., the Central Solenoid Converter Power Supply grid [49]) coupled with real-time simulation strategies to monitor and study the performance of the ITER grid while the reactor is functional.
Distributed models and dynamic data
@@
For coupling simulations across multiple HPC centers, it is possible to rely on a single workspace that provides a global view of information shared from multiple geo-distributed HPC data centers [53]. The feasibility of distributed workflows has been shown by coupling the Rosetta-based scientific tools with interactive notebook-based analyses executed outside the HPC center [54]. Another application of this approach is for neutron scattering science [55], for which the performance challenges related to the transfer of large files between the HPC center and the experimental facility have been identified as one of the main performance bottlenecks for the deployment of workflow motif. A complementary work to these studies investigates the overhead and performance of coupling non-HPC systems (equivalent to edge devices) to handle HPC workflows [56].
@@
(a) HPC simulations coupled with experiments and AI surrogate models and AI analytics geographically distributed using dynamic interaction and resource management
(b) Adaptive methods for training ensemble of models with large datasets on HPC potentially coupled with simulations that generate the data needed for training

### v2

@@ Abstract
AI integration is revolutionizing the landscape of HPC simulations, enhancing the importance, use, and performance of AI-coupled HPC workflows. This paper surveys the diverse and rapidly evolving field of AI-driven HPC and provides a common conceptual basis for understanding such workflows. Specifically, we use insights from different modes of coupling AI into HPC workflows to propose six execution motifs most commonly found in scientific applications. By definition, the proposed set of execution motifs is incomplete and evolving. However, they allow us to analyze the primary performance challenges across AI-integrated simulation environments. We close with a listing of open challenges, research issues, and suggested areas of investigation, including the the need for specific benchmarks that will help evaluate and improve the execution of AI-coupled HPC workflows.
Keywords: scientific workflows, high performance computing, machine learning, deep learning, active learning
@@ 1 Introduction
Multiple recent publications have demonstrated how artificial intelligence (AI) coupled with traditional high performance computing (HPC) simulations can provide practical performance enhancements of 103 or more [1, 2, 3]. AI-coupled HPC workflows – as opposed to decoupled AI and HPC – involve the concurrent, real-time coupled execution of AI and HPC tasks in ways that allow the AI systems to steer or inform the HPC tasks and vice versa. The real-time coupling and concurrency of HPC and AI tasks are fundamental, as they allow bidirectional influence. The online coupling of an AI system to HPC workflows can be used for a variety of scenarios, e.g., an AI-surrogate model to substitute part of expensive simulation, an AI system to learn the function or determine the parameters or the practical fields, and an AI system to guide a campaign using an objective function or performance Pareto-optimal or resource-optimal experiments (using, for example, Active/Reinforcement Learning).
@@
Driving the need to couple AI systems to traditional HPC simulations is the promise of enhancing the “effective performance” of HPC workflows, where enhancement is measured by “science for a given amount of computing”. Such coupled workflows continue to overcome the limitations of traditional forward simulations in increasingly sophisticated and pervasive ways. Further, these hybrid workflows will overcome traditional bottlenecks that prevent greater scale – physical and temporal or higher resolutions. The integration of AI into a computational workflow is also a sustainable and scalable way to obtain significant performance gains and present an opportunity to avoid simulation enhancements that are overly sensitive to processor architecture [4, 5]. Put together, AI-coupled HPC workflows present a promising paradigm [6, 7, 8] which leverages the ubiquitous interest and wide capabilities being developed for AI but employs them to overcome performance bottlenecks due to unsustainable approaches.
Integrating AI systems into traditional high-performance computing workflows has demonstrably enabled highly accurate modeling and holds significant promise for accelerating scientific discovery. There exist multiple promising examples of workflows that couple AI methods with traditional HPC workflows, typically by extending HPC workflows with additional capabilities to support the concurrent execution of AI modules  [9, 10, 11, 12]. Multiple solutions have successfully coupled AI systems to be either “about” the HPC workflows or “outside” (typically in the outer loop or even possibly remote from) the main HPC workflow. The integration of AI with HPC can be systematically understood through a taxonomy that defines three key modes of interaction: AI-in-HPC, AI-out-HPC, and AI-about-HPC [6]. AI-in-HPC signifies instances where an AI model directly replaces or acts as a surrogate for an entire HPC simulation or one of its components, thereby accelerating computations. In contrast, AI-out-HPC describes scenarios where an AI model operates externally to the core HPC simulation loop but dynamically controls or orchestrates the progression of the HPC workflow, such as through active learning or reinforcement learning to optimize computational campaigns. Finally, AI-about-HPC refers to situations where AI models function concurrently and are coupled with the primary HPC tasks, serving to complement traditional computational methods by enhancing scientific results or improving overall efficiency without necessarily taking over core simulation functions.
Figure 1 shows a typical science application where AI has advanced simulations and experiments by coupling AI to HPC workflows via different mechanisms.
Hybrid AI-HPC applications offer significant improvements over traditional high-performance computing (HPC) methods by fundamentally reshaping parallel computation, resource utilization, and scientific insight extraction. Unlike vanilla HPC applications, which rely heavily on exhaustive, synchronous computations across HPC platforms, AI-integrated workflows employ predictive inference, intelligent execution, and optimization to achieve scientific outcomes.
Traditional HPC methods necessitate step-by-step simulation runs, which can be expensive to explore even modest design spaces. In contrast, AI-guided simulations leverage trained AI models to predict system behaviors, significantly reducing computational burdens. AI models, after training on extensive datasets from high-fidelity simulations, can almost instantaneously forecast outcomes, bypassing the exhaustive computations previously required. While the initial AI model training might demand significant computational resources, subsequent inferences are orders of magnitude faster and more cost-effective. This enables real-time optimizations, rapid design-space exploration, and immediate predictions, thereby dramatically streamlining workflows and accelerating scientific discoveries.
Furthermore, AI frameworks naturally employ asynchronous computation, fundamentally differing from traditional HPC’s synchronous, lockstep patterns. AI workloads, which can tolerate and even thrive on asynchronous updates, such as certain neural network training algorithms, minimize synchronization overheads, a critical bottleneck in large-scale HPC scenarios. This asynchronous execution allows different parts of computations to progress independently, significantly improving efficiency by not tying the entire system to the slowest processing elements.
Integrating AI into HPC workflows restructures parallel algorithm designs and communication patterns, directly addressing and overcoming the strong scaling limits inherent to traditional HPC simulations. AI intelligently predicts outcomes, identifies critical computation regions, and streamlines workflows by learning underlying physical laws directly from data or encoding known physics into AI models. These predictive and adaptive capabilities minimize communication overhead, balance computational loads, and accelerate computationally intensive segments without sacrificing accuracy, thereby pushing beyond traditional scaling limitations and significantly enhancing scalability.
AI-coupled HPC workflows support algorithms/methods where machine learning complements traditional solvers by initializing simulations, preconditioning algorithms, or estimating error bounds, enhancing both computational efficiency and scientific reliability. Building on this integration, surrogate modeling allows AI to approximate complex simulations, delivering results significantly faster than conventional approaches. Through accelerated simulations, AI focuses computational resources on the most promising regions of parameter space using techniques like active learning, thereby reducing cost and time. Hybrid approaches as exemplified above, yield faster scientific insights, broader design explorations, representing an essential shift away from traditional performance gains and limitations of strong scaling.
@@
The focus in this paper is on the scientific workflows that involve the online (real-time) coupling and concurrent execution of AI and HPC, not offline training and inference. In addition to providing an overview of existing AI-coupled HPC workflows, the primary objectives of this paper are to identify classes of AI-coupled HPC workflows and to provide an overview of the software systems used to support them. The paper also discusses performance and systems challenges, identifying robust solutions and lacking ones. Finally, the paper highlights some open research issues to advance AI-HPC workflows.
The paper is organized as follows: Section 2 reviews coupling modes between AI and HPC tasks and introduces six execution motifs that describe the interaction and coordination patterns of AI and HPC components. These motifs are used to understand a range of AI-coupled HPC workflow applications and their associated performance challenges. Execution motifs are important at multiple levels, not the least of which is their influence on the design of application-level software systems and task runtime systems. For instance, the frequency of interaction in a given motif informs the performance requirements and the optimizations available to lower-level libraries. Further, the diversity of AI approaches and model types mean intrinsic variation in training and execution times. Section 3 presents several frameworks and libraries developed to ease the coupling of AI and HPC components in a single workflow and illustrates how each of these frameworks relates to the proposed high-level modes of execution motifs. Section 4 discusses various performance issues related to AI-coupled HPC workflows – organized around the motifs – such as load balancing, workload management, and dataflow performance bottlenecks. We conclude in Section 5 with a discussion of several open issues and challenges that this emerging class of workflows presents.
@@ 2 Execution Motifs
The concept of AI-coupled HPC workflows is well-established, with a current taxonomy dividing them into three primary modes: AI-in-HPC, AI-out-HPC, and AI-about-HPC [6, 5]. These modes are characterized by the relative placement of AI methods concerning HPC simulations. AI-in-HPC represents the scenario in which an AI system is introduced instead of a component of the HPC simulation, or possibly, instead of the whole simulation itself, i.e., the AI model serves as a “total surrogate” [13]. AI-out-HPC captures situations where an AI system resides “outside” of the traditional HPC simulation loop but dynamically controls the progression of the HPC workflow [14, 15] (e.g., control of computational campaigns via reinforcement learning). Finally, AI-about-HPC represents the situation where AI systems are concurrent and coupled to the main HPC tasks: analysis and training codes use the output of the HPC simulation to provide further insights. This section expands on this taxonomy. Given the increasing trend and benefits of combining these non-mutually exclusive modes [4, 16, 5], the need for compatible software solutions for managing their execution and data exchange is paramount. For this purpose, we build on the concept of coupling modes (AI-x-HPC, where x could be in, out, or about), and we introduce execution motifs. Whereas modes provide insight into the static coupling of AI systems with HPC simulations, execution motifs capture and provide information on recurring dynamic interaction patterns between the AI systems and the HPC simulations, driven by a high-level functional goal.
@@
These execution motifs serve as identifiers for (groups of) AI-coupled HPC workflows that exhibit similar interaction and coupling patterns between AI and HPC components, and thus encounter similar performance bottlenecks and require similar optimizations when deployed on HPC systems. Because a workflow might incorporate aspects of several motifs, grasping the relationships between workflow patterns, the frameworks employed, and the performance considerations outlined in this paper empowers users to better comprehend the available options and choose appropriate solutions tailored to their multi-faceted workflow needs. Figure 2 illustrates the characteristics used to differentiate between different workflow patterns and how they can be combined to create an example motif.
@@
Figure 2: Defining characteristics of AI-coupled HPC workflows: (left) Breakdown of the main characteristics used to define the different behavioral motifs; (right) Illustrative example of the characteristics typically used in the steering motif following the breakdown presented on the left.
Based on a literature survey of large-scale workflows, we identify six recurring execution motifs that capture the dominant coupling patterns between AI and HPC. These motifs reflect the dynamic interaction styles and purposes of AI within scientific workflows. Table 1 summarizes the distinct patterns of each motif according to the characteristics introduced in Figure 2. We briefly list them here:
  1. 1.
    Dynamic Orchestration: AI steers ensembles of HPC simulations, dynamically spawning or terminating runs.
  2. 2.
    Multistage Pipeline: AI-based filters guide execution across stages in a pipelined HPC campaign.
  3. 3.
    Inverse Design: AI iteratively infers simulation inputs from observations to solve inverse problems.
  4. 4.
    Digital Replica: AI digital replicas interact with simulations for prediction, monitoring, and feedback.
  5. 5.
    Distributed Models: AI and HPC components operate across geographically distributed resources.
  6. 6.
    Adaptive Training: HPC simulations generate data that adaptively improve AI models during training.
Figure 3: Sankey diagram of the motifs and their characteristics.
Motif Patterns Example Use Case
Dynamic
Orchestration
@@
[Uncaptioned image] - Command-and-control of physical experiments and simulations (e.g. between shots feedback for plasma physics)
@@
Multistage Pipeline [Uncaptioned image] - Large-scale MD simulations using AI sampling of a system with many degrees of freedom
@@
InverseDesign [Uncaptioned image] - Materials discovery to address the problem of data sparsity and reduce the need for domain-specific knowledge
@@
Digital Replica [Uncaptioned image] - Digital twin of a fusion reactor running concurrent with ITER [17] and monitoring
@@
Distributed Models [Uncaptioned image] - Edge-to-HPC execution of large-scale simulations
- Interactive notebook-based analyses of instrument data and AI output
AdaptiveTraining [Uncaptioned image] - Hyperparameter optimization workflows
@@
Dynamic Orchestration
@@
((a)) AI system steering (ensembles of) high-fidelity HPC simulations to avoid undesired results or guide the workflow towards better solutions.
((b)) Sequential coupling of data-dependent multistage AI-coupled HPC workflow as a pipeline with filtering/analysis codes between stages.
Figure 4: Workflow computational patterns for the Dynamic Orchestration (a) and Multistage Pipeline (b) motifs
The following literature survey focuses on workflows for the Dynamic Orchestration execution motif. These workflows use AI systems to analyze how “well” (ensembles of) simulations are progressing toward their initial scientific objective. Based on that analysis, they can change the “trajectory” when exploring the parameter space. This allows them to avoid undesired results either or increase performance or accuracy. The training of the AI system is done offline before the execution of the workflow. Workflows in which the steering process is also used to guide the online training of the AI system will be investigated as part of the Adaptive Training motif.
@@
DeepDriveMD [2] is a general-purpose framework to run an analysis code (as a separate application) that pre-processes the simulation data (e.g. selecting only a subset of atoms of interest or calculating physical parameters for a protein’s native state), and then uses machine learning (ML) models on the pre-processed data to decide which simulations to run next and terminates the less productive Molecular Dynamics (MD) simulations. By focusing on promising regions of the parameter space, AI-based steering can efficiently navigate complex, high-dimensional systems that were previously intractable. AI systems can also significantly reduce the computational resources required to obtain accurate results by steering simulations toward more relevant configurations.
The aforementioned example is a typical implementation of the Dynamic Orchestration motif, but other examples using different coupling solutions while sharing the characteristics presented in Table 1 exist. The AI system analyzes the output of the HPC simulation and either steers the existing simulation or triggers new simulations based on the outcome.
@@
High-Throughput Virtual Screening Pipelines are emblematic of this motif. For instance, a virtual drug discovery pipeline can use AI to improve the effective sampling of individual stages [24]. Integrating multiple AI-coupled workflows into a single multistage pipeline is another example of AI-driven discovery [25]. AI can also be used in workflows composing multiple multi-fidelity simulations to dynamically adjust how many and what fidelity levels to run based on the desired outcome [26].
Less complex multistage pipelines can be found in neuroscience [27, 28] and cancer research [29]. Both fields employ workflows composed of a sequence of medical image processing applications that process the same Magnetic Resonance Image (MRI) or Whole Slide Image (WSI) with the output of one application informing the following decision. Each image-processing application is typically a mix of HPC and AI codes composed of multiple stages, including cleaning, standardizing, analysis, and post-processing within one large code or as separate applications. One such workflow is, for instance, used to identify lesions in the brain [28]. It consists of a 3-stage pipeline, starting with a set of MRI images that go through a whole segmentation code, followed by a surface reconstruction application, and finally by an application to create a structure of the brain in MRI space. Based on the output of these three applications, the workflow can further trigger additional codes depending on the research objectives.
@@
Deep learning has been used to solve inverse problems, i.e., determining causal factors from a set of observations, in many scientific fields, including material science [32, 33], nanophotonics [34, 35, 36], or heat transfer [37]. However, the lack of sufficiently large training data sets is often a crucial issue for these learning algorithms since their accuracy heavily relies on the amount and the quality of the available data. This is, for instance, the case with material characteristics identification, which requires a huge amount of user time to acquire observational data. HPC simulations have recently started generating sufficient noise-free data to train AI models. The new challenge is getting accurate simulation results, often requiring extensive testing and expertise to fine-tune the simulation input parameters. This can also be time-consuming and resource-intensive, especially for complex simulations since it requires validation over many tests. AI-coupled HPC workflows following the Inverse Design motif address this data sparsity problem and reduce the demand for domain-specific knowledge in determining the input parameters of a simulation.
@@
((a)) Iterative refinement of design parameters combining experiments with HPC simulations and AI surrogates
((b)) Physics-based models combining HPC high-fidelity simulations and AI reduced order models and/or digital twins
@@
Fusion energy researchers are turning to digital twins as a promising avenue for efficient whole-device optimization and virtual component qualification, crucial for successfully designing fusion reactor components under conditions that defy real-world testing. Digital twin workflows are primed to play a crucial role in ensuring the success of the international thermonuclear experimental reactor (ITER) [17], and pave the way for the future of commercial fusion energy. Several studies [48, 49, 50] presents the vision of the fusion science community to develop future workflows. They all discuss the design of a workflow combining digital twins of different parts of the fusion reactor (e.g., the Central Solenoid Converter Power Supply grid [48]) coupled with real-time simulation strategies to monitor and study the performance of the ITER grid while the reactor is functional.
Distributed models
@@
For coupling simulations across multiple HPC centers, it is possible to rely on a single workspace that provides a global view of information shared from multiple geo-distributed HPC data centers [52]. The feasibility of distributed workflows has been shown by coupling the scientific tools for macromolecular modeling and protein design with interactive notebook-based analyses executed outside the HPC center [53]. Another application of this approach is for neutron scattering science [54], for which the performance challenges related to the transfer of large files between the HPC center and the experimental facility have been identified as one of the main performance bottlenecks for the deployment of workflow motif. A complementary work to these studies investigates the overhead and performance of coupling non-HPC systems (equivalent to edge devices) to handle HPC workflows [55].
@@
((a)) HPC simulations coupled with experiments and AI surrogate models and AI analytics geographically distributed using dynamic interaction and resource management
((b)) Adaptive methods for training ensemble of models with large datasets on HPC potentially coupled with simulations that generate the data needed for training